Calculations: For a mass of 250 grams hung at 0 degrees and another 250 gram mass hung at 90 degrees: Experiment 1: A) What is the magnitude and direction of the net, unbalanced force for this situation?
Answer #1

A) By Newtons law of Gravitation,

F = Gm1m2/r^2

= 6.67e-11*0.25*0.25/r^2

= 4.17*10^-12/r^2

Here r is distance between the masses whoch would have been given in the experiment.

Here F is so small with any r, that we can ignore them.

So the only significant forces will be Tension and weight mg.

Weight mg =0.25*9.8 = 2.45N acts downwards always.